OpenStudy (anonymous):

Solve for the variable: m/4.2 =5 A. 25/21 B. -25/21 C. -21 D. 21

OpenStudy (saifoo.khan):

\[m = 5*4.2\]

OpenStudy (anonymous):

(m/4.2)= 5/1 and cross multiply to solve ans isolate x which gives you D. 21

OpenStudy (anonymous):

Mantra: repeat over and over, "what you do on one side you must do on the other."

OpenStudy (saifoo.khan):

i mean't multiply. Sorry

OpenStudy (anonymous):

thank you.